Circular No. 6820

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1998O, 1998P, 1998Q L. Germany reports that the Mount Stromlo Abell Cluster Supernova

Search Team (cf. IAUC 6639) has discovered three apparent supernovae on V and R CCD images (limiting mag V = 21.5) taken by R. Hola at the Mount Stromlo 1.27-m telescope (+ Macho Camera):

SN  1998 UT  α_{2000}  δ_{2000}  V  Offset
1998O Jan. 27 13^{h}39^{m}34.^{s}98 −35^{o}46′32′′.5 20.9 1′′.1 E, 16′′.4 N
1998P Jan. 27 13 37 17.63 −35 27 32.6 20.1 2′′.3 E, 2′′.5 N
1998Q Jan. 27 13 32 46.38 −31 37 23.5 18.5 0′′.1 E, 1′′.5 N

SNe 1998O, 1998P, and 1998Q are located in anonymous galaxies near the Abell galaxy clusters A3566, A3566, and A3562, respectively. Confirmation images (limiting mag V = 21.5) for SN 1998Q were obtained by S. Sabine with the same telescope on Jan. 31 and for the supernovae in A3566 on Feb. 4. None of the supernovae were present on images taken on 1997 July 17 (limiting mag V = 20). Finding charts are available at

http://msowww.anu.edu.au/~reiss/Abell SNSearch/.

SUPERNOVA 1997ef IN UGC 4107 L. Wang, D. A. Howell, and J. C. Wheeler, University of Texas at

Austin, report: “SN 1997ef was observed on 1997 Dec. 28 and 1998 Feb. 2 UT with the 2.1-m telescope at McDonald Observatory. The 1998 Feb. data show that SN 1997ef is making the transition to the nebular phase, and the spectrum strongly resembles those of type-Ic supernovae at 1–2 months past maximum (similar to the type-Ic spectrum of SN 1997ei taken on the same night). Broad emission lines of He I (587.6 nm)/Na I (589.0, 589.6 nm), [O I] (630.0, 636.4 nm), O I (777.1 nm), and Mg I] (457.1 nm) were detected. Spectropolarimetry was also obtained, but no polarization was detected down to ∼ 1.5 percent.”

COMET 103P/HARTLEY 2 Total visual magnitude estimates: Jan. 17.73 UT, 8.4 (M. Meyer,

Frauenstein, Germany, 20×100 binoculars); 21.12, 8.0 (C. E. Spratt, Victoria, BC, 0.08-m refractor); 25.74, 8.7 (M. Reszelski, Szamotuly, Poland, 20×60 binoculars); 27.81, 8.6 (A. Pereira, Cabo da Roca, Portugal, 14×100 binoculars); 31.99, 8.8 (J. Bortle, Stormville, NY, 20×80 binoculars); Feb. 3.08, 9.8 (G. W. Kronk, Troy, IL, 0.33-m reflector; moonlight).

1998 February 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green